OpenStudy (anonymous):

Solve the system, using substitution. 4x + y = 4 and y = 4 - 4x A. no solution B. infinitely many solutions

hartnn (hartnn):

plug in y= 4-4x in your first equation 4x +y =4 what do u get?

hartnn (hartnn):

means replace 'y' in 4x+y = 4 by 4-4x

OpenStudy (anonymous):

have you tried to do what hartnn suggested ?

OpenStudy (anonymous):

B. Infinite solutions?

hartnn (hartnn):

yes! how did u get it ?

OpenStudy (anonymous):

By plugging in.

hartnn (hartnn):

ok, thats correct :)